# A Low-Rank ADMM Splitting Approach for Semidefinite Programming

### 14 Mar 2024 v1

@@ 7      Set $k\leftarrow k+1$
$\max\left\{\frac{\|\mathcal{A}(U^{k+1}(V^{k+1})^{\top})-b\|}{1+\|b\|_{\infty}},\frac{\gamma\|V^{k+1}-V^{k}\|_{F}}{1+\|V^{k}\|_{F}}\right\}<\epsilon$ (13)
@@
Assume $\|V^{k}\|_{F}\leq\sigma$ and $(U^{k}+1,V^{k+1},\lambda^{k+1})\in\left(\mathcal{B}_{\delta}\big{(}0^{n\times r}\big{)}\right)^{2}\times\mathcal{B}_{\delta}\big{(}0^{m}\big{)}$. By using triangle inequality, it is not difficult to show
@@
This implies that the $k$-th iterate $(U^{k+1},V^{k+1},\lambda^{k+1})$ is a $\max\left\{c,1+\|b\|_{\infty}\right\}\epsilon$-KKT point of problem (6) and thus $(\frac{U^{k+1}+V^{k+1}}{2},2\lambda^{k+1})$ is an $O(\epsilon)$-KKT point of (4) by taking $\gamma=O(1/\sqrt{\epsilon})$.
@@ 2.3 The Power of ADMM
The intuition for constructing the bilinear factorization stems from the observation that within the iterative trajectory of LRADMM, $U$ and $V$ do not need to be identical. This splitting affords a more flexible search area, potentially endowing the algorithm with the advantage of concurring hard instances. Figure 1 depicts a scenario in which the Burer-Monteiro approach is unable to solve the SDP relaxation of a sensor network localization (SNL) problem, whereas LRADMM demonstrates success. Given the location of anchors observed distances information between sensor-sensor and sensor-anchor pairs, the objective of SNL is to determine the sensor’s true position, represented by the blue star. In a more formal setting, for a $d$-dimensional graph $G(V,E)$, we differentiate between the sensor set $V_{x}=\{x_{1},x_{2},\cdots,x_{n}\}$ and the anchor set $V_{a}=\{a_{1},a_{2},\cdots,a_{m}\}$ being a partition of $V$, where these sets collectively partition $V$ and $m\geq d+1$. Observable edges are represented by two sets: $N_{x}={(i,j)\mid i,j\in V_{x}}$ for sensor-sensor pairs and $N_{a}={(i,j)\mid i\in V_{a},j\in V_{x}}$ for sensor-anchor pairs. The lengths of these edges are denoted as $d_{ij}=\left\|x_{i}-x_{j}\right\|,\leavevmode\nobreak\ \forall(i,j)\in N_{x}$ for sensor-sensor edges and $\tilde{d}_{kj}=\left\|a_{k}-x_{j}\right\|,\leavevmode\nobreak\ \forall(k,j)\in N_{a}$ for sensor-anchor edges. The objective of the SNL problem is formulated as follows:
$\min_{x}\sum_{\{i,j\}\in N_{x}}\left|\left\|x_{i}-x_{j}\right\|^{2}-d_{ij}^{2}\right|+\sum_{(k,j)\in N_{a}}\left|\left\|a_{k}-x_{j}\right\|^{2}-\tilde{d}_{kj}^{2}\right|,$
@@
$Z=\left[\begin{array}[]{cc}I&X\\ X^{T}&Y\end{array}\right]$
@@ 3 Convergence Analysis of LRADMM
In this section, we present the main theoretical results of our proposed Algorithm 1. We will prove the convergence of ADMM for solving the bilinear factorized problem and further improve the convergence rate under the assumption of Łojasiewicz inequality.
@@ 3.1 Descent Property of LRADMM
Suppose that Assumption 1 holds. Let $\big{\{}U^{k},V^{k},\lambda^{k}\big{\}}_{k=0}^{K-1}$ be the sequence generated by Algorithm 1, then for any $k\geq 0$, we have
$\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k+1})-\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k})=\tfrac{1}{\rho}\|{\lambda}^{k+1}-{\lambda}^{k}\|^{2},$ (17)
$\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k})-\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\leq-\tfrac{\gamma}{2}(\|V^{k+1}-V^{k}\|^{2}+\|U^{k+1}-U^{k}\|^{2}).$ (18)
It is worth noting that the update of the dual variable ${\lambda}$ leads to an increase in the value of the augmented Lagrangian function. Therefore, bounding $\|\lambda^{k+1}-\lambda^{k}\|^{2}$ is crucial to ensure a sufficient decrease in each iteration. To this end, we need some boundedness and nonsingularity conditions.
@@
There exists a $\rho_{0}>0$ such that for any given $\beta$, the level set $S_{\beta}:=\{(U,V)\mid f(UV^{\top})+\frac{\rho_{0}}{2}\|U-V\|_{F}^{2}+\frac{\rho_{0}}{2}\|{\cal A}(UV^{\top})-b\|^{2}\leq\beta\}$ is either compact or empty.
Assumption 3.
There exists a constant $\eta>0$ such that
$\min_{(U,V)\in S_{\hat{\beta}}}\max_{R\in\mathrm{O}_{n}}\sigma_{\min}\big{(}C(UR)\big{)}\geq\eta,$
where $\mathrm{O}_{n}$ represents the set of orthogonal matrices of dimension $n$, function $C:\mathbb{R}^{n\times r}\rightarrow\mathbb{R}^{nr\times m}$ is defined as $C(U)=[\operatorname{vec}{(}A_{1}U),\cdots,\operatorname{vec}{(}A_{m}U)]$, and $\sigma_{\min}$ denotes the least singular value of a matrix.
@@
    $f(X)+\frac{\rho_{0}}{2}\|\mathcal{A}(X)-b\|^{2}$ strongly convex with respect to $X\in\mathbb{R}^{n\times n}$.
@@
$\mathcal{L}_{\rho}(U^{0},V^{0},{\lambda}^{0})+\frac{1}{\rho}\|{\lambda}^{0}\|^{2}\leq\beta^{0}<\hat{\beta},$ (19)
@@
With these foundations in place, we are now prepared to establish the bound of $\|\lambda^{k+1}-\lambda^{k}\|^{2}$ and hence the descent property of our proposed Algorithm 1.
@@
    $\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\leq\beta^{0}$ and $\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})+\frac{3}{4\rho}\|{\lambda}^{k}\|^{2}<\hat{\beta}$;
@@
    $\|{\lambda}^{k+1}-{\lambda}^{k}\|^{2}\leq c_{1}\|\Delta U^{k+1}\|^{2}+c_{2}\|\Delta V^{k+1}\|^{2},$ where $c_{1}=\frac{2}{\eta^{2}}\big{(}L^{\delta}(\delta)^{2}+L_{f}^{\delta}+\gamma+\frac{1}{\eta}L_{f}^{\delta}\delta\sqrt{s_{A}}+\frac{2}{\eta}\gamma\delta\sqrt{s_{A}}\big{)}^{2}$ and $c_{2}=\frac{2}{\eta^{2}}\big{(}L^{\delta}\delta^{2}+\gamma\big{)}^{2}$;
@@
    $\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k+1})-\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\leq-(\frac{\gamma}{2}-\frac{c_{1}}{\rho})\|\Delta U^{k+1}\|^{2}-(\frac{\gamma}{2}-\frac{c_{2}}{\rho})\|\Delta V^{k+1}\|^{2}.$
@@ 3.2 Convergence and Convergence Rate of ADMM
The augmented Lagrangian function $\mathcal{L}_{\rho}(U,V,{\lambda})$ satisfies the Łojasiewicz inequality with exponent $\alpha\in[0,1)$ at any KKT point of problem (6), i.e. there exist constants $c,\delta_{c}>0$ such that for any $(U,V,{\lambda})\in\mathbb{R}^{n\times r}$ and $(\overline{U},\overline{V},\overline{\lambda})\in\mathcal{W}$ satisfying $\operatorname{dist}\big{(}(U,V,{\lambda}),(\overline{U},\overline{V},\overline{\lambda})\big{)}\leq\delta_{c}$,
@@
Suppose Assumptions 1– 3 hold. Let the initial point $(U^{0},V^{0},{\lambda}^{0})$ and $\rho,\gamma$ satisfy the conditions of Lemma 1. Then, the sequence of iterates $\{(U^{k},V^{k},{\lambda}^{k})\}_{k\geq 0}^{K-1}$ converges to a KKT point of problem (6), $(\overline{U},\overline{V},\overline{{\lambda}})$. Moreover, the total iteration complexity to achieve an $\epsilon$-KKT point is $\mathcal{O}(\epsilon^{-1})$.
By precisely characterizing the exponent $\alpha$ of the Łojasiewicz inequality, we can determine the convergence rate of the ADMM.
@@ 4.1 LoRADS
LoRADS is designed to derive a satisfactory low-rank approximate solution for the linear SDP problem (2). The approach unfolds in two distinct phases: Phase I centers on the Burer-Monteiro factorization problem (4). To solve this problem, we employ the augmented Lagrangian algorithm (Burer and Monteiro, 2003), initiating with a random starting point and employing the L-BFGS method to enhance algorithmic efficacy.
In the ensuing discussion, we refer to the augmented Lagrangian algorithm applied to solve problem (4) as LRALM. When problem (4) is solved to meet certain accuracy, we stop the augmented Lagrangian algorithm and take the secured solution $(\tilde{U}^{*},\tilde{{\lambda}^{*}})$. Then we switch to Phase II, focusing on solving the penalized bilinear decomposed problem (6) with LRADMM. We split the matrix variable $U$ and initiate LRADMM by setting $U^{0}=V^{0}=\tilde{U}^{*},{\lambda}^{0}=\tilde{\frac{1}{2}{\lambda}^{*}}$, where $(U^{0},V^{0},{\lambda}^{0})$ is expected to be a satisfactory initial point as stated in Lemma 1. The LRADMM is then executed to derive an approximate solution $(\overline{U},\overline{V})$ for problem (6) satisfying the final tolerance. By setting $\widehat{U}=\frac{\overline{U}+\overline{V}}{2}$, we obtain an intermediate solution for problem (4). Finally, with $\widehat{X}=\widehat{U}\widehat{U}^{\top}$, we get a solution to the general SDP problem (1). The complete pseudocode for the LoRADS is presented in Algorithm 2.
@@ 4.2 The Motivation of the Two-phases Structure
Although we aim at developing an approach based on the previously proposed LRADMM to tackle SDP problems, our exploration reveals that directly applying Algorithm 1 to solve (6) may not guarantee stability for all classes of general SDPs. Through empirical analysis, we sometimes observe poor performance when initiating from randomly selected points. Notwithstanding, Theorem 2 indicates that Algorithm 1 exhibits promising convergence properties, if the initial point $(U^{0},V^{0},{\lambda}^{0})$ aligns with the conditions specified in Lemma 1. These conditions essentially suggest that $U^{0}$ and $V^{0}$ are situated within an area near the optimal solutions $U^{*}$ and $V^{*}$. In light of this, we propose a warm-start approach. Naturally, we use the Burer-Monteiro method (Burer and Monteiro, 2003) for warm-start, as it keeps the low-rank factorization and directly yields a factorized solution $X=UU^{\top}$. To illustrate the ideas discussed above, we compare the warm-start approach against applying LRADMM throughout and applying LRALM throughout on a set of representative problems selected [^1] They are selected from the problems tested in Section 5.. In this experiment, we maintain a constant rank of $\sqrt{2m}$ and other employ SDPLR as the solver for LRALM. Other implementation details and experiment settings follow Section 5.1 and 5.2.
@@
The outcomes, presented in Table 1, enable us to make several key observations: 1) ADMM demonstrates its edge against the ALM approach on problems such as G60, G64 and theta12; 2) in contrast, for problems like MC_3000 and H3O, ADMM without a warm-start is notably slower or may even fail; 3) the warm-started LRADMM shows a stable performance across the problems tested. Therefore, to enhance ADMM’s stability, we integrate a warm-start phase using LRALM.
It is important to remark, however, that the algorithm for solving problem (4) in Phase I is not exclusively limited to LRALM. The LoRADS approach is designed to be compatible with various methods used for warm-start in Phase I. More empirical results presented in Section 5 indicate that using LRALM for the warm-start phase is sufficiently effective, though exploring more effective warm-start strategies remains a potential avenue for future investigation.

### 25 Mar 2024 v2

@@ 7      Set $k\leftarrow k+1$
$\max\left\{\frac{\|\mathcal{A}(U^{k+1}(V^{k+1})^{\top})-b\|_{2}}{1+\|b\|_{\infty}},\frac{\gamma\|V^{k+1}-V^{k}\|_{F}}{1+\|V^{k}\|_{F}}\right\}<\epsilon$ (13)
@@
Assume $\|V^{k}\|_{F}\leq\sigma$ and $(U^{k+1},V^{k+1},\lambda^{k+1})\in\left(\mathcal{B}_{\sigma}\big{(}0^{n\times r}\big{)}\right)^{2}\times\mathcal{B}_{\sigma}\big{(}0^{m}\big{)}$. By taking the Frobenius norm on both sides of (14) and using triangle inequality, it is not difficult to show
@@
This implies that the $k$-th iterate $(U^{k+1},V^{k+1},\lambda^{k+1})$ is a $(\max\left\{c,1+\|b\|_{\infty}\right\}\cdot\epsilon)$-KKT point of problem (6) and thus $\big{(}\frac{U^{k+1}+V^{k+1}}{2},2\lambda^{k+1}\big{)}$ is an $O(\epsilon)$-KKT point of (4) by taking $\gamma=\Omega(1/\sqrt{\epsilon})$.
@@ 2.3 The Power of ADMM
The intuition for constructing the bilinear factorization stems from the observation that within the iterative trajectory of LRADMM, $U$ and $V$ do not need to be identical. This splitting affords an enlarged search area, which sometimes improves the algorithm performance on hard instances. Figure 1 depicts a scenario in which the Burer-Monteiro approach is unable to solve the SDP relaxation of a sensor network localization (SNL) problem, whereas LRADMM demonstrates success. Given the location of anchors observed distances information between sensor-sensor and sensor-anchor pairs, the goal of SNL is to determine the sensor’s true position, represented by the blue star. In a more formal setting, for a $d$-dimensional graph $G(V,E)$, we differentiate between the sensor set $V_{x}=\{x_{1},x_{2},\cdots,x_{n}\}$ and the anchor set $V_{a}=\{a_{1},a_{2},\cdots,a_{m}\}$ being a partition of $V$, where these sets collectively partition $V$ and $m\geq d+1$. Observable edges are represented by two sets: $N_{x}=\{(i,j)\mid i,j\in V_{x}\}$ for sensor-sensor pairs and $N_{a}=\{(i,j)\mid i\in V_{a},j\in V_{x}\}$ for sensor-anchor pairs. The lengths of these edges are denoted as $d_{ij}=\left\|x_{i}-x_{j}\right\|_{2},\leavevmode\nobreak\ \forall(i,j)\in N_{x}$ for sensor-sensor edges and $\tilde{d}_{kj}=\left\|a_{k}-x_{j}\right\|_{2},\leavevmode\nobreak\ \forall(k,j)\in N_{a}$ for sensor-anchor edges. The SNL problem is formulated as follows:
$\min_{x}\sum_{(i,j)\in N_{x}}\left|\left\|x_{i}-x_{j}\right\|_{2}^{2}-d_{ij}^{2}\right|+\sum_{(k,j)\in N_{a}}\left|\left\|a_{k}-x_{j}\right\|_{2}^{2}-\tilde{d}_{kj}^{2}\right|,$
@@
$Z=\left[\begin{array}[]{cc}I&X\\ X^{\top}&Y\end{array}\right].$
@@ 3 Convergence Analysis of LRADMM
In this section, we present the main theoretical results of our proposed Algorithm 1. We will prove the convergence of LRADMM for solving the bilinear factorized problem and further improve the convergence rate under the assumption of Łojasiewicz inequality.
@@ 3.1 Descent Property of LRADMM
Suppose that Assumption 1 holds. Let $\big{\{}(U^{k},V^{k},\lambda^{k})\big{\}}_{k=0}^{K-1}$ be the sequence generated by Algorithm 1, then for any $k\geq 0$, we have
$\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k+1})-\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k})=\tfrac{1}{\rho}\|{\lambda}^{k+1}-{\lambda}^{k}\|_{2}^{2},$ (17)
$\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k})-\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\leq-\tfrac{\gamma}{2}(\|V^{k+1}-V^{k}\|_{F}^{2}+\|U^{k+1}-U^{k}\|_{F}^{2}).$ (18)
It is worth noting that the update of the dual variable ${\lambda}$ leads to an increase in the value of the augmented Lagrangian function. Therefore, bounding $\|\lambda^{k+1}-\lambda^{k}\|_{2}^{2}$ is crucial to ensure a sufficient decrease in each iteration. To this end, we need some boundedness and nonsingularity conditions.
@@
There exists a $\rho_{0}>0$ such that for any given $\beta$, the level set $S_{\beta}:=\{(U,V)\mid f(UV^{\top})+\frac{\rho_{0}}{2}\|U-V\|_{F}^{2}+\frac{\rho_{0}}{2}\|{\cal A}(UV^{\top})-b\|_{2}^{2}\leq\beta\}$ is either compact or empty.
@@
    $f(X)+\frac{\rho_{0}}{2}\|\mathcal{A}(X)-b\|_{2}^{2}$ strongly convex with respect to $X\in\mathbb{R}^{n\times n}$.
@@
$\mathcal{L}_{\rho}(U^{0},V^{0},{\lambda}^{0})+\frac{1}{\rho}\|{\lambda}^{0}\|_{2}^{2}\leq\beta^{0}<\hat{\beta},$ (19)
@@
We also need the following assumption, which assumes that the constraints $\mathcal{A}(UU^{T})=b$ are linearly independent within the region $S_{\hat{\beta}}$. More previously, let the function $\mathcal{C}:\mathbb{R}^{n\times r}\rightarrow\mathbb{R}^{nr\times m}$ denote the map from any point $U$ to the matrix consisting of the vectorized gradients of these constraints $\mathcal{A}(UU^{T})=b$, i.e., $\mathcal{C}(U)=[\operatorname{vec}{(}A_{1}U),\cdots,\operatorname{vec}{(}A_{m}U)]$. We assume that $\mathcal{C}(U)$ is of full rank for any point $(U,V)$ in the region $S_{\hat{\beta}}$.
Assumption 3.
There exists a constant $\eta>0$ such that
$\min_{(U,V)\in S_{\hat{\beta}}}\sigma_{\min}\big{(}\mathcal{C}(U)\big{)}\geq\eta,$
where and $\sigma_{\min}$ denotes the least singular value of a matrix.
Assumption 3 is called the linear independence constraint qualification (LICQ), which is standard in the convergence analysis of ADMM-type methods. For example, see Lin et al. (2015); Wen et al. (2010); Xie and Wright (2021).
With these foundations in place, we are now prepared to establish the bound of $\|\lambda^{k+1}-\lambda^{k}\|_{2}^{2}$ and hence the descent property of our proposed Algorithm 1.
@@
    $\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\leq\beta^{0}$ and $\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})+\frac{3}{4\rho}\|{\lambda}^{k}\|_{2}^{2}<\hat{\beta}$;
@@
    $\|{\lambda}^{k+1}-{\lambda}^{k}\|_{2}^{2}\leq c_{1}\|\Delta U^{k+1}\|_{F}^{2}+c_{2}\|\Delta V^{k+1}\|_{F}^{2},$ where $c_{1}=\frac{2}{\eta^{2}}\big{(}L^{\delta}\delta^{2}+L_{f}^{\delta}+\gamma+\frac{1}{\eta}L_{f}^{\delta}\delta\sqrt{s_{A}}+\frac{2}{\eta}\gamma\delta\sqrt{s_{A}}\big{)}^{2}$ and $c_{2}=\frac{2}{\eta^{2}}\big{(}L^{\delta}\delta^{2}+\gamma\big{)}^{2}$;
@@
    $\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k+1})-\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\leq-(\frac{\gamma}{2}-\frac{c_{1}}{\rho})\|\Delta U^{k+1}\|_{F}^{2}-(\frac{\gamma}{2}-\frac{c_{2}}{\rho})\|\Delta V^{k+1}\|_{F}^{2}.$
@@ 3.2 Convergence and Convergence Rate of ADMM
The augmented Lagrangian function $\mathcal{L}_{\rho}(U,V,{\lambda})$ satisfies the Łojasiewicz inequality with exponent $\alpha\in[0,1)$ at any KKT point of problem (6), i.e. there exist constants $c,\delta_{c}>0$ such that for any $(U,V,{\lambda})\in\mathbb{R}^{n\times r}\times\mathbb{R}^{n\times r}\times\mathbb{R}^{m}$ and $(\overline{U},\overline{V},\overline{\lambda})\in\mathcal{W}$ satisfying $\operatorname{dist}\big{(}(U,V,{\lambda}),(\overline{U},\overline{V},\overline{\lambda})\big{)}\leq\delta_{c}$,
@@
Suppose Assumptions 1– 4 hold. Let the initial point $(U^{0},V^{0},{\lambda}^{0})$ and $\rho,\gamma$ satisfy the conditions of Lemma 1. Then, the sequence of iterates $\{(U^{k},V^{k},{\lambda}^{k})\}_{k\geq 0}^{K-1}$ converges to a KKT point of problem (6), $(\overline{U},\overline{V},\overline{{\lambda}})$. Moreover, the total iteration complexity to achieve an $\epsilon$-KKT point is $\mathcal{O}(\epsilon^{-1})$.
Therefore, we have established the $O(1/k)$ convergence rate of our ADMM method. However, as observed in our numerical experiments, Algorithm 1 shows an even faster convergence rate in most cases. This is achievable if we could precisely characterize the exponent $\alpha$ of the Łojasiewicz inequality.
@@ 4.1 LoRADS
LoRADS is designed to derive a satisfactory low-rank approximate solution for the linear SDP problem (2). The approach unfolds in two phases: Phase I centers on the Burer-Monteiro factorization problem (4). To solve this problem, we employ the augmented Lagrangian algorithm (Burer and Monteiro, 2003), initiating with a random starting point and employing the L-BFGS method to enhance algorithmic efficacy.
In the ensuing discussion, we refer to the augmented Lagrangian algorithm applied to solve problem (4) as LRALM. When problem (4) is solved to a certain accuracy, we stop the augmented Lagrangian algorithm and take the secured solution $(\tilde{U}^{*},\tilde{\lambda}^{*})$. Then, we switch to Phase II, focusing on solving the penalized bilinear decomposed problem (6) with LRADMM. We split the matrix variable $U$ and initiate LRADMM by setting $U^{0}=V^{0}=\tilde{U}^{*},{\lambda}^{0}=\frac{1}{2}\tilde{\lambda}^{*}$, where $(U^{0},V^{0},{\lambda}^{0})$ is expected to be a satisfactory initial point as stated in Lemma 1. The LRADMM is then executed to derive an approximate solution $(\overline{U},\overline{V})$ for problem (6) satisfying the final tolerance. By setting $\widehat{U}=\frac{\overline{U}+\overline{V}}{2}$, we obtain an intermediate solution for problem (4). Finally, with $\widehat{X}=\widehat{U}\widehat{U}^{\top}$, we get a solution to the general SDP problem (1). The complete pseudocode for the LoRADS is presented in Algorithm 2.
@@ 4.2 The Motivation of the Two-phases Structure
Although we aim at developing an approach based on the previously proposed LRADMM, our exploration reveals that directly applying Algorithm 1 to solve (6) might not always ensure stability. Empirically, initiating from arbitrary starting points occasionally leads to suboptimal performance. In view of our theoretical result (Theorem 2), Algorithm 1 exhibits promising convergence properties when the initial point $(U^{0},V^{0},{\lambda}^{0})$ aligns with the conditions specified in Lemma 1. These conditions imply that $U^{0}$ and $V^{0}$ should be close to the optimal solutions $U^{*}$ and $V^{*}$ to ensure fast convergence. To address this, we propose a warm-start approach, which initializes ADMM by the Burer-Monteiro method (Burer and Monteiro, 2003), as BM keeps the low-rank factorization and directly yields a factorized solution $X=UU^{\top}$. To illustrate the ideas discussed above, we compare the warm-start approach against applying LRADMM throughout and applying LRALM throughout on a set of representative problems selected [^1] They are selected from the problems tested in Section 5.. In this experiment, we maintain a constant rank of $\sqrt{2m}$ and employ SDPLR as the solver for LRALM. Other implementation details and experiment settings follow Section 5.1 and 5.2.
@@
  • 2
    For LRADMM, we use an initial $\rho=100$.
From Table 1, we observe that while LRADMM outperforms the LRALM in solving problems such as G60, G40_mb and theta12, it can be notably slower and may even fail in the other cases. However, the warm-started LRADMM exhibits a stable performance across the problems tested. Therefore, to enhance ADMM’s stability, we integrate a warm-start phase using LRALM. More empirical results presented in Section 5 indicate that using LRALM for the warm-start phase is adequately effective, though exploring more effective warm-start strategies presents an intriguing direction for future research.